ROTATIONAL FINE STRUCTURE IN THE FLUORESCENCE EXCITATION SPECTRUM OF TOLUENE

\underline{ALLAN L. L. EAST}, I. DECHENE, W. SIEBRAND, M. Z. ZGIERSKI, P. R. BUNKER, Steacie Institute of Molecular Science, National Research Council, Ottawa, Ontario K1A 0R6, Canada; P. JENSEN, FB 9-Theoretische Chemie, Bergische Universitat-Gesamthochschule Wuppertal, D-42097 Wuppertal, Germany.

Experimental measurements have been reported the S_{1}−S_{0} fluorescence excitation spectrum of toluene^{1} which reveal forbidden bands corresponding to excited levels of the methyl internal rotation mode. We present theoretical predictions of the rotational fine structure of these bands, incorporating ab initio and least-squares-fitted geometrical structures for the S_{0} and S_{1} states. We particularly consider the coupling of overall and internal rotation, and the direct calculation of individual rovibronic line intensities.

1 R. A. Walker, E. Richard, K.-T. Lu, E. L. Sibert III, and J. C. Weisshaar, J. Chem. Phys. 102 (1995) p. 8718, and references therein.